OpenStudy (anonymous):

The distance between the two parallel lines L1: −7 x − 10 y=−4, L2: −7 x − 10 y=3 is ??????

OpenStudy (sriram):

7

OpenStudy (anonymous):

no its not 7

OpenStudy (anonymous):

L1: −7 x − 10 y=−4, L2: −7 x − 10 y=3 when x1=1 , y=-1.1 also x2=1, y=-1 using the distance formula d=sqrt[(x1-x2)^2+(y1-y2)^2] d=sqrt[(1-1)^2 +(-1.1+1)^2 d=sqrt[0+(-0.1)^2] d=sqrt[0.1=0.1 ans]

OpenStudy (anonymous):

L1: −7 x − 10 y=−4, L2: −7 x − 10 y=3 when x1=1 , y=-3/10 when x2=1, y=-1 using the distance formula d=sqrt[(x1-x2)^2+(y1-y2)^2] d=sqrt[(1-1)^2 +(-3/10-(-1))^2 d=sqrt[0+(-3/10-(-1))^2] d=sqrt[((-3/10)+1)^2] d=sqrt[((-3+10)/10)^2] d=sqrt[(7/10)^2] d=7/10 d=0.7